OpenStudy (shamil98):

Evaluate the integral: \[\large \int\limits_{-1}^{-8} \frac{ x-x^2 }{ 2 \sqrt[3]{x} }\]

OpenStudy (anonymous):

It's just integrating power functions.

OpenStudy (shamil98):

something like this right? \[= \frac{ x }{ 2(x)^{\frac{ 1 }{ 3 }} } - \frac{ x^2 }{ x^\frac{ 1 }{ 3 } }\] \[= \frac{ \frac{ x^2 }{ 2*x^{\frac{ 4 }{ 3 }} } } {\frac{ 4 }{ 3 } } - \frac{\frac{ x^3 }{ 3x^{\frac{ 4 }{ 3} } }}{\frac{ 4 }{ 3 }}\]

OpenStudy (shamil98):

Actually, I just noticed I could simplify it first..

OpenStudy (anonymous):

something like \[ \frac{x}{\sqrt[3]{x}} = x^{1-1/3} = x^{2/3} \]

OpenStudy (shamil98):

yeah, overthinking this simplified to this, and then integrate it.. \[\large \frac{ x^{2/3} }{ 2 } - \frac{ x^{5/3} }{ 2 }\]

OpenStudy (anonymous):

Yes

OpenStudy (shamil98):

\[= \frac{ 5x^{5/3} }{ 6 } - \frac{ 8x^{8/3} }{ 6 }\]

OpenStudy (shamil98):

\[=\frac{ 5(-8)^{5/3} }{ 6 } - \frac{ 8(-8)^{8/3} }{ 6 } -\left( \frac{ 5(-1)^{5/3} }{ 6 } - \frac{ 8(-1)^{8/3} }{ 6 } \right)\] ?

OpenStudy (shamil98):

I think my calculus teacher purposely gives these problems a lengthy amount of work -.-

OpenStudy (shamil98):

is this right so far?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

I like how your calc teacher makes you integrate backwards.

OpenStudy (shamil98):

He's ancient.

OpenStudy (shamil98):

I think i made an error with my integration though..

OpenStudy (shamil98):

\[\int\limits_{}^{} \frac{ x^{2/3} }{ 2 } = \frac{ 3x^{5/3} }{ 10 }\] right?

OpenStudy (nincompoop):

is your last post the problem?

OpenStudy (shamil98):

no, the problem is the definite integral, i'm just doing part of the integration again because i think it was done incorrectly..

OpenStudy (nincompoop):

integral of [x^(2/3)]2 dx?

OpenStudy (nincompoop):

oh okay

OpenStudy (shamil98):

yeah, i checked with wolfram, that part was wrong. so uh: \[\huge \int\limits_{-8}^{-1} \frac{ (x-x^2) }{ 2\sqrt[3]{x} } =>\frac{ x }{ 2\sqrt[3]{x} } - \frac{ x^2 }{ 2\sqrt [3]{x} } =\] \[\huge \frac{ x^{2/3} }{ 2 } - \frac{ x^{5/3} }{ 2 }\] \[\huge = \frac{ 3x^{5/3} }{ 10 } - \frac{ 3x^{8/3} }{ 16 }\]

OpenStudy (shamil98):

\[\large = \frac{ 3(-1)^{5/3} }{ 10 } - \frac{ 3(-1)^{8/3} }{ 16 } - \left( \frac{ 3(-8)^{5/3} }{ 10 } - \frac{ 3(-8)^{8/3} }{ 16 } \right)\] evaluating this part is going to be a pain..

OpenStudy (nincompoop):

how so?

OpenStudy (shamil98):

i'm a lazy person, thats how

OpenStudy (nincompoop):

good enough reason come up with a technique to make it easier

OpenStudy (shamil98):

i'll leave that to the geniuses